Mar 31, 2010 Calorimetry

From GlueXWiki
Jump to: navigation, search

Teleconference Time: 11:00 a.m. EST

  • ESNET and EVO session
  • Phone in CC F326 is 757-269-6337
  • Phone in the Regina Videoconference Suite is 306-585-4204
  • Phone only if requested:
    • +1-800-377-8846 : US
    • +1-888-276-7715 : Canada
    • +1-302-709-8424 : International
    • then enter participant code: 77438230# (remember the "#"). Phone line by request only.

Items for followup from previous meeting(s)

  1. Monitoring
    • Matt to coordinate with Christina on the shipment of glass blocks and pmts/dividers for monitoring system test.
    • Bcal: Testing to see if LED sources can be viewed by pmts on both sides of the module (UofR).
  2. Fcal
    • Decision on rad hard detectors
    • Check instability of FEU-84 pmt gains / finalize number / order additional FEU-84 pmts PR for 450 pmts issued

Documents to Review

Tentative Agenda

  1. Announcements
  2. Review of Action Items
  3. Calibration/Monitoring
  4. FCAL
    • Construction status
    • Readiness Review
    • To shashlyk or not to shashlyk
    • Resolution of the Radiation Hard Glass -- File:RHGupdate.pdf
  5. BCAL
    • Construction Status
  6. Simulations
  7. Other Updates


Attendees: Fernando, Pedro, John, Elton, Beni, Tim, Eugene (JLab); Mat, Mijahlo, Claire, Brad, Ryan (IU); George, Andrei (UofR).

  1. Announcements
    1. Today is actually April 1st!
    2. Zisis had to make an unplanned trip back to Greece to take care of his mother. Dan will be on vacation next week.
  2. Review of Action Items
    1. Matt will contact Christina on shipment of lead glass
    2. Bcal testing
      • George reported that the setup is proceeding well. Jigs have been machined (90% complete) to attach light guides at each end. Will check normalize left and right with cosmic rays. Tests of LED sources will follow
    3. Elton has submitted a PR for the order of 450 FEU-84-3 pmts.
      • Matt inquired about the testing of the pmts. If schedule matches, could be a good project for summer student.
  3. Calibration/Monitoring
  4. FCAL
    1. Construction status
      • Drawings have been set out to get quotes for materials
      • Exploring how best to do the accounting for the project
      • Assembly technician advertised; should interview in 2 weeks
      • Paul, John and Matt visited Diversified Systems, a nearby electronic firm to check facilities. Paul was very impressed. The company would be able to go from drawings to complete assembly including manual soldering. Comparing costs for complete assembly compared to board fabrication/stuffing with completion at IU.
    2. Readiness Review
      • Tentative schedule is first week in May.
    3. To shashlyk or not to shashlyk
    4. Resolution of the Radiation Hard Glass -- File:RHGupdate.pdf
      • Comparisons of simulations using HDGeant and GEANT4 model. In particular, the measured properties of rad-hard F108 glass have been added to the program (previous simulations used the nominal properties of F101). Goal is to update the HDGeant model so that it reproduces the measured properties of rad-hard glass.
  5. BCAL
    1. Construction Status
      • Saturday, modules 1-4 were checked for transmission. Good uniformity (few percent) measured. Photos will be made available.
      • Precision of holes for pins along the length were very good.
      • Measurement of vertical position of holes difficult using the corner as a reference. However, the precision is considered sufficient for our specifications. Future measurements will use a jig that allows an easier reference than the corner.
      • Modules will be made ready for shipment once measurements are complete
      • Modules 5 and 6 matrices are complete, will soon be transported to Ross
      • Modules 7 and 8 stacking has started
      • Documentation for shipment of 1-4 is being finalized
      • All other production operations are normal.
      • Uniformity studies: Uniformity meets specifications, but variations are thought to be due to the rubber. Small modifications for improvement are tested. These include the removal of the silicon rubber and use of thinner polyethylene, which seems to result in better uniformity.
  6. Simulations
    1. 2.5 GeV momenta at 12deg was started a few days ago; should be completed soon. Then the photoelectron stats and spectra will be made.
    2. Next on the list is the threshold (60 MeV) upstream.
    3. Finally a run at 2.5 GeV at 14deg will be run to compare maximum energies with the 12deg run.
  7. Other Updates